OpenStudy (toxicsugar22):

Dichloro is a pesticide that was used in the middle decades of the 20th century to control maleria. Afterr 1945, it was also widely used on crops in the U, and as much as one ton be sprayed on a single cotton field. Hoowever after the toxic, effects of Dichloro on the evvironment began to appear, the chemical was banned in 1972. A common estimate for the half-life in the soil is 15 years. Write a decay law for Dichloro in the soil and use A for the initial concentration of Dicloro, and use t for the input variable.

OpenStudy (dumbcow):

\[y = Ae ^{-kt}\] \[\frac{1}{2}A=Ae ^{-15k}\] \[\frac{1}{2} = e ^{-15k}\rightarrow 2 = e ^{15k}\] \[k = \frac{\ln2}{15}\]

OpenStudy (toxicsugar22):

that the answer

OpenStudy (dumbcow):

yes well you have to plug that k back into the top equation

OpenStudy (toxicsugar22):

so it is\[y=Ae-1n2/15^t\]

OpenStudy (toxicsugar22):

is that right

OpenStudy (dumbcow):

correct :)

OpenStudy (toxicsugar22):

\[y=Ae-1n2/15^t\]

OpenStudy (dumbcow):

is that a subtractions sign ??? \[y = Ae ^{-\frac{\ln 2}{15}t} \]

OpenStudy (toxicsugar22):

THANKS